OpenStudy (anonymous):

complete the table of values. write the results as ordered pairs. 7x-3y=21 the first ordered pair, where x=0,_is the second ordered pair is where y=0,_ is the third ordered pair is where y= -14,_is

OpenStudy (nikita2):

x=0 -3y=21 => y = -7 y=0 7x=21 => x =3 y=-14 7x - 3*14 = 21 x = 9

OpenStudy (anonymous):

x=0, y= -7 y=0, x =3 y= -14 , x= -3